# Exploring Exponential Runge-Kutta Methods: A Survey

### v2

@@ 5. Examples
To conclude our manuscript, we present three examples useful for demonstrating the functionalities of order 2 ExpRK method (ExpRK2) and their comparison with other numerical methods of equal order, namely the standard order 2 RK method (RK2) and the classic order 2 Rosenbrock method (Rb2). In this way, we aim to provide both a more complete description of the numerical methods analyzed in this manuscript and a useful tool for approaching ExpRK methods. We remark that, in 2020, Montanelli and Bootland [montanelli2020solving] published a comprehensive review of exponential methods, comparing various formulas for solving stiff PDEs in multiple dimensions. Their analysis, focused on periodic boundary conditions and semilinear stiff PDEs, concluded that the ETDRK4 scheme by Cox and Matthews [cox2002exponential] (denoted as ExpRK4 in the following, for ease and uniformity of notation) remains one of the most effective approaches despite the proliferation of more complex alternatives. Their work is particularly notable for its examination of both diffusive problems (characterized by large negative eigenvalues) and dispersive problems (distinguished by their highly oscillatory nature), as well as its discussion of computational techniques for the $\varphi$-functions based on contour integrals. In line with Montanelli and Bootland’s review [montanelli2020solving], we restrict our attention to explicit exponential methods. These methods are by far the most widely employed in practice: they often yield superior performance when stiffness is predominantly carried by the linear dynamics while the nonlinear contribution remains comparatively mild. Moreover, they avoid the extra complications associated with Newton-type iterations, such as the need to solve nonlinear systems at every step and to tune iterative solver tolerances; these steps bring additional computational cost and potential convergence issues. For these reasons, explicit exponential formulations are a natural and pragmatic choice in many applications. Nevertheless, in this work we also deliberately test them on less typical, small-scale examples to probe their behavior in unconventional settings.
@@
The three didactic examples we propose have dimension 1, 2 and 50 (where here the ‘dimension’ refers to the size of the system of ODEs), respectively. Specifically, for the 1D case, we revisit the introductory example presented by Cox and Matthews in 2002 [cox2002exponential, Sect. 4.1]; for the 2D case, we examine a classical model from nonlinear mechanics and dynamical systems, namely the Duffing oscillator [^12] For a comprehensive analysis of the nonlinear Duffing oscillator and its applications, we recommend the following volume [DuffingBook] to the interested reader.; the last example is the result of applying the pseudospectral discretization method [bdgsv16] to an Ikeda DDE [ELLG04]. In what follows, we briefly introduce three order $2$ methods and three order $4$ methods for solving stiff systems of ODEs. As already noticed, stiff ODEs require specialized numerical techniques due to their rapidly changing components existing alongside slowly evolving ones. The six methods under consideration–RK2, ExpRK2, Rb2, RK4, ExpRK4 and Rb4–represent different approaches to handling stiffness, while all maintaining second-order accuracy. Their properties suggest different performance characteristics when applied to stiff problems.
@@
Among the order 4 exponential methods, we consider ExpRK4 [cox2002exponential, eq. (26)-(29)]. Considering again the equation,
@@
for constant $\alpha_{ij},\,\beta_{i},\,\gamma_{ij}$ which are chosen in order to satisfy the order 4 conditions (see, e.g., [hw91, Section IV.7]).
@@
Following the same approach presented by Cox and Matthews [cox2002exponential] in their analysis of ETD methods, we consider a non-autonomous ODE model with rapid decay, characteristic of stiff systems:
@@
The behavior of this differential equation can be divided into two distinct phases: an initial fast transient phase, during which the solution rapidly approaches the slow manifold, followed by a phase in which the solution evolves along this slow manifold. As noticed in [cox2002exponential], when $k<0$ and $|k|\gg 1$, the solution quickly approaches a slow manifold on which
@@
These findings align with the conclusions drawn by Cox and Matthews in [cox2002exponential], confirming the superiority of exponential methods for solving stiff problems. In particular, ExpRK2 offers a significant advantage in terms of stability and accuracy, permitting the use of much larger step sizes than conventional methods. The superiority of ExpRK4 with respect to the tested methods of the same order is even more evident. For stiff problems such as the one studied here, characterized by a parameter k=100k=-100, exponential methods represent the optimal choice among the three methods of the corresponding order, striking an ideal balance between accuracy, stability, and computational efficiency.
@@
The results presented in Figure 2 and Tables 4-5 provide a comprehensive analysis of six numerical methods of order 2 and 4 applied to the Duffing oscillator system [^13] All relative errors are computed with respect to a high-precision numerical solution obtained using MATLAB’s ode15s solver, as no analytical solution is available for this nonlinear system. with $\omega=1$ and $k=100$.
@@
Lastly, we consider the Ikeda DDE, which models a variety of passive optical systems subject to a delayed feedback [ELLG04]. The general form of the DDE reads
@@
In order to make (35) amenable to our analysis, we need to first discretize the DDE into a finite number $M$ of ODEs. We opt in particular for the pseudospectral discretization technique [bdgsv16]. In fact, the ODEs obtained using this approach are defined by a semilinear right-hand side with stiff linear part, regardless of the right-hand side of the DDE. The resulting non-linear term is defined by only one nonzero component, which corresponds to the non-linear term in the original right-hand side (see [ando2024proceedings] for details on this approach).
For our experiments, we choose $M=50$, $\lambda=40$ and $\mu=2$. The dimension of the resulting ODE is thus much larger than the previous examples, yet smaller than many applications which involve the discretization of a PDE. The linear part is dominant but, again contrary to the majority of examples analyzed in the literature of exponential methods, it is not sparse. We compute the error with respect to a reference solution obtained with dde23, a MATLAB integrator for DDEs. The final error, therefore, includes both the contribution of the discretization and that of the numerical method for approximating the solution.
@@ 6. Conclusions
The dual nature of this contribution, at once a scholarly survey and a didactic introduction, represents, we believe, its distinctive value. On the one hand, specialists may find here a structured synthesis of the major milestones, including recent advances such as high stiff-order methods, parallelization strategies, and applications in modeling contexts. On the other hand, students and non-experts may use the text as a stepping stone, moving from the basic principles to more sophisticated topics without losing sight of the underlying intuition. The coexistence of these two layers reflects the current state of the field itself, which remains both historically rich and dynamically evolving.
@@ Acknowledgements
AA is member and acknowledges the support of Gruppo Nazionale per il Calcolo Scientifico (GNCS) of Istituto Nazionale di Alta Matematica (INdAM).

### v3

@@ 5. Examples
To conclude our manuscript, we present three examples that illustrate the general functionality of ExpRK methods. To this end, we consider standard methods of two different orders (2 and 4) and compare an ExpRK method with a Rosenbrock method and a classical explicit RK method of the same order. The purpose of these examples is not to draw general conclusions in favor of exponential methods, nor to test specific features or provide an exhaustive assessment of their performance. Rather, our goal is to gain insight into which classes of equation could benefit from the use of ExpRK methods, by means of comparison with well-established classical schemes. In this perspective, the examples are meant to have a primarily illustrative and didactic value, offering a practical introduction to the behavior of ExpRK methods in representative settings.
@@
We remark that, in 2020, Montanelli and Bootland [83] published a comprehensive review of exponential methods, comparing various formulas for solving stiff PDEs in multiple dimensions. Their analysis, focused on periodic boundary conditions and semilinear stiff PDEs, concluded that the ETDRK4 scheme by Cox and Matthews [20] (denoted as ExpRK4 in the following, for ease and uniformity of notation) remains one of the most effective approaches despite the proliferation of more complex alternatives. Their work is particularly notable for its examination of both diffusive problems (characterized by large negative eigenvalues) and dispersive problems (distinguished by their highly oscillatory nature), as well as its discussion of computational techniques for the $\varphi$-functions based on contour integrals. In line with Montanelli and Bootland’s review [83], we restrict our attention to explicit exponential methods. These methods are by far the most widely employed in practice: they often yield superior performance when stiffness is predominantly carried by the linear dynamics while the nonlinear contribution remains comparatively mild. Moreover, they avoid the extra complications associated with Newton-type iterations, such as the need to solve nonlinear systems at every step and to tune iterative solver tolerances; these steps bring additional computational cost and potential convergence issues. For these reasons, explicit exponential formulations are a natural and pragmatic choice in many applications. Nevertheless, in this work we also deliberately test them on less typical, small-scale examples to probe their behavior in unconventional settings.
@@
The three didactic examples we propose have dimension 1, 2 and 50 (where here the ‘dimension’ refers to the size of the system of ODEs), respectively. Specifically, for the 1D case, we revisit the introductory example presented by Cox and Matthews in 2002 [20, Sect. 4.1]; for the 2D case, we examine a classical model from nonlinear mechanics and dynamical systems, namely the Duffing oscillator [^12] For a comprehensive analysis of the nonlinear Duffing oscillator and its applications, we recommend the following volume [59] to the interested reader.; the last example is the result of applying the pseudospectral discretization method [11] to an Ikeda DDE [33]. In what follows, we briefly introduce three order $2$ methods and three order $4$ methods for solving stiff systems of ODEs. As already noticed, stiff ODEs require specialized numerical techniques due to their rapidly changing components existing alongside slowly evolving ones. The six methods under consideration represent different approaches to handling stiffness, while all maintaining second-order [resp. fourth-order] accuracy. Their properties suggest different performance characteristics when applied to stiff problems.
@@
Among the order 4 exponential methods, we consider ExpRK4 [20, eq. (26)-(29)]. Considering again the equation,
@@
for constant $\alpha_{ij},\,\beta_{i},\,\gamma_{ij}$ which are chosen in order to satisfy the order 4 conditions (see, e.g., [40, Section IV.7]).
@@
Following the same approach presented by Cox and Matthews [20] in their analysis of ETD methods, we consider a non-autonomous ODE model with rapid decay, characteristic of stiff systems:
@@
The behavior of this differential equation can be divided into two distinct phases: an initial fast transient phase, during which the solution rapidly approaches the slow manifold, followed by a phase in which the solution evolves along this slow manifold. As noticed in [20], when $k<0$ and $|k|\gg 1$, the solution quickly approaches a slow manifold on which
@@
These findings align with the conclusions drawn by Cox and Matthews in [20], confirming the superiority of exponential methods for solving stiff problems. In particular, ExpRK2 offers a significant advantage in terms of stability and accuracy, permitting the use of much larger step sizes than conventional methods. The superiority of ExpRK4 with respect to the tested methods of the same order is even more evident. For stiff problems such as the one studied here, characterized by a parameter k=100k=-100, exponential methods represent the optimal choice among the three methods of the corresponding order, striking an ideal balance between accuracy, stability, and computational efficiency.
@@
The results presented in Figure 2 and Tables 4-5 provide a comprehensive analysis of six numerical methods of order 2 and 4 applied to the Duffing oscillator system [^13] All relative errors are computed with respect to a high-precision numerical solution obtained using MATLAB’s ode15s solver, as no analytical solution is available for this nonlinear system. integrated over the time interval $[0,20]$ with $\omega=1$ and $k=100$.
@@
Lastly, we consider the Ikeda DDE, which models a variety of passive optical systems subject to a delayed feedback [33]. The general form of the DDE reads
@@
In order to make (35) amenable to our analysis, we need to first discretize the DDE into a finite number $M$ of ODEs. We opt in particular for the pseudospectral discretization technique [11]. In fact, the ODEs obtained using this approach are defined by a semilinear right-hand side with stiff linear part, regardless of the right-hand side of the DDE. The resulting non-linear term is defined by only one nonzero component, which corresponds to the non-linear term in the original right-hand side (see [3] for details on this approach).
For our experiments, we chose $M=50$, $\lambda=40$ and $\mu=2$, and integrated the equation over the time interval $[0,4]$. The dimension of the resulting ODE is thus much larger than the previous examples, yet smaller than many applications which involve the discretization of a PDE. The linear part is dominant but, again contrary to the majority of examples analyzed in the literature of exponential methods, it is not sparse. We compute the error with respect to a reference solution obtained with dde23, a MATLAB integrator for DDEs. The final error, therefore, includes both the contribution of the discretization and that of the numerical method for approximating the solution.
@@ 6. Conclusions
The dual nature of this contribution, at once a scholarly survey and a didactic introduction, represents, we believe, its distinctive value. On the one hand, specialists may find here a structured synthesis of the major milestones, including recent advances such as high stiff order methods, parallelization strategies, and applications in modeling contexts. On the other hand, students and non-experts may use the text as a stepping stone, moving from the basic principles to more sophisticated topics without losing sight of the underlying intuition. The coexistence of these two layers reflects the current state of the field itself, which remains both historically rich and dynamically evolving.
@@ Acknowledgements
AA is member and acknowledges the support of Gruppo Nazionale per il Calcolo Scientifico (GNCS) of Istituto Nazionale di Alta Matematica (INdAM). Her work was partially supported by the Italian Ministry of University and Research (MUR) through the PRIN 2022 project (No. 20229P2HEA) “Stochastic numerical modelling for sustainable innovation”, Unit of Udine (CUP G53C24000710006).